# SoK: Membership Inference Attacks on LLMs are Rushing Nowhere (and How to Fix It)

### v2

@@ 6.4 Post-hoc control methods
Sequence The randomized split of The Pile [31], with additional deduplication, as made available by Duan et al. [27] in the MIMIR dataset [^5] https://huggingface.co/datasets/iamgroot42/mimir. We recommend using the split where sequences with more than 80% of $13$-gram overlapping are not removed, not considering data from Github. Suite of Pythia [7] and GPT-NEO [8] models. Allows for randomized MIA evaluation. In this setup, MIAs are more challenging due to the inherent overlap between members and non-members This however represent a more realistic setup in general.
@@
Sequence The randomized unique traps sequences as made available by Meeus et al. [70], across all number of repetitions $(10,100,1,000)$. CroissantLLM [30] Allows for randomized MIA evaluation in a realistic real-world model, limiting the overlap between members and non-members, and across different memorization regimes. MIA performance might not be realistic, and might be specific to sequence characteristics such as length, and perplexity.
@@
Document Randomized set of ArXiv papers sampled from the train and test splits from The Pile [31], both available as complete documents as well as in sequences. Suite of Pythia [7] and GPT-NEO [8] models. Allows for randomized MIA evaluation. In this setup, MIAs are more challenging due to the inherent overlap between members and non-members This however represent a more realistic setup in general.
@@
Document Set of ArXiv papers sampled close to the training data cut-off date for OpenLLaMA[32]. Suite of OpenLLaMA[32] models. Allows for MIA evaluation on data collected post-hoc with any distribution shift between members and non-members largely mitigated (RDD). Might still be subject to a distribution shift, and faces the overlap issue.
Table 6: Recommended benchmarks to evaluate new MIAs against LLMs proposed in the future, highlighting benefits and limitations.
6.4.1 Beating the bag of words baseline
A potential solution is to revisit the baseline used to infer binary membership. MIAs against machine learning (ML) models traditionally use a 50/50 split of membership for evaluation, resulting in a random guess AUC of $0.5$ [88, 12, 18, 57, 84, 96, 35, 53, 75, 82, 106]. Alternatively, Jayaraman et al. [42] argues that a balanced prior for membership is not realistic in practice and proposes to evaluate MIAs with skewed priors. More recently, Kazmi et al. [47] proposes privacy auditing of ML models using a baseline from a (model-less) classifier.
Hence, we could set the performance of the bag of words classifier as a prior. An effective MIA methodology should then outperform this baseline by leveraging access to the target model and exploiting its memorization of member target records.
This solution has significant limitations in practice. First, the performances of the bag of words classifier applied to the widely used datasets for post-hoc MIAs (see Table 2) are near-perfect. With such a significant distribution shift between members and non-members, it would become practically impossible for novel MIAs to surpass this baseline and might lead us to (incorrectly) conclude that the target model does not memorize. Second, such model-less baselines will always be a lower bound. In Sec. 4, we only used a simple bag-of-words classifier, since its performance was already substantial. Had this classifier performed closer to random guessing, a subtler distribution shift might still exist, potentially detectable by more sophisticated classifiers [111, 24]. Hence, unless the distribution shift is strongly mitigated, using a model-less prior is not a valid option.
6.4.2 Regression Discontinuity Design (RDD)
@@
In the domain of causal inference, Regression Discontinuity Design (RDD) is a technique used to study the effect of a treatment when the assignment to the treatment has not been randomized but is instead based on a cutoff point. By sampling observations in a principled manner close to this cutoff, RDD enables the estimation of the treatment effect even in the absence of a randomized setup. RDD is for instance widely used in econometrics [50], where cutoff points might include income levels for welfare programs or geographical borders when studying different policy implications.
@@
We instantiate this RDD setup on OpenLLaMA [32], a model for which the entire training dataset is known. As members, we consider papers from a month before the cutoff date for the training dataset of the target LLM (February ’23). For non-members, we consider papers released in March ’23 — so one month after the member documents.
@@
Table 8 (Appendix B) summarizes the AUC for both sequence- as well as document-level MIAs across the literature applied to this setup. We find that the performance of MIAs remains close to a random guess baseline, as also found in the case of a randomized test split 6.1.
@@
While the RDD method we propose seems to offer a viable solution for the setup at hand (using OpenLLaMa [32] and ArXiv papers), its practical implementation may face limitations. First, the RDD setup assumes knowledge of the training data cut-off date (which is likely quite different from the model release date) and relies on the availability of sufficient data released close to this date. For ArXiv papers, more than $10,000$ new papers are added every month  [^6] https://arxiv.org/stats/monthly_submissions, allowing for ample data collection close to the cut-off date. While similar arguments can be made for other frequently updated data sources (e.g. news articles), the RDD approach could prove challenging for less frequently updated sources (e.g. crime novels). Second, the RDD method aims to significantly mitigate distribution shifts between members and non-members. The quantification of such biases faces similar issues as discussed above. Lastly, the RDD setup does not address the issue of potential overlap between members and non-members. This might be preferred if one wants to calibrate realistic MIA performance in practice, but maybe less so when benchmarking new methods.
6.4.3 Debiasing techniques
@@
Recently, Eichler et al.[29] introduced two novel methods to mitigate bias between members and non-members collected post-hoc. The first method creates “non-biased” datasets, where members and non-members share similar n-gram distributions. The second one creates “non-classifiable” datasets, which resist baselines not taking the target model into account such as the bag of words classifier we here propose. While authors show these techniques to eliminate biases present in the original datasets collected post-hoc, this approach faces similar limitations as other post-hoc control methods. Fully eliminating known biases, not introducing new biases and accurately evaluating their presence, remains a practical challenge.
7 Conclusion
@@
In this work, we identified that recent research on MIAs against LLMs often uses flawed evaluation setups. In the last year, research has proposed at least $10$ different MIA methodologies, whose performance has been impossible to fairly and appropriately compare and benchmark.
After carefully examining potential solutions, and their drawbacks, we identify a set of benchmarks to use when new MIAs are proposed. We summarize our findings in Table 6.
@@
For sequence-level MIAs, we propose a combination of the MIMIR dataset [27] (with limited deduplication) and the randomized trap sequences provided by Meeus et al. [70]. MIMIR enables a randomized setup for MIA evaluation across different model sizes [7, 8] and, despite its overlap issues, serves as a realistic benchmark for MIAs in practice. The trap sequences allow for benchmarking new MIA methodologies across different levels of memorization regimes; however, the corresponding results may depend on the specific properties of these sequences.
@@
Finally, we encourage future research on MIAs against LLMs to always include a model-less baseline and to remain mindful of the inherent challenges (e.g. overlap, scaling from finetuning to pretraining) before interpreting results.
Availability. The code and data required to replicate the results in this paper, as well as to benchmark new MIAs against LLMs are available on Github.
Appendix A Detailed MIA implementation
Table 7 elaborates on the implementation details used for the MIAs from the literature we consider. Note that some MIAs, i.e. Loss [106], Zlib [16] and Lower [16], do not depend on the choice of certain hyperparameters.
@@ Appendix A Detailed MIA implementation
PAC [105] Using the implementation provided by Ye et al. [105], we generate $10$ augmentations with alpha parameter of $0.3$. We the compute polarized distance using $k_{m}ax=0.05$ and $k_{m}in=0.3$ following the best performing results reported in the original paper.
@@
Table 7: MIA Implementation Details Summary
Note that we do not consider the MIAs requiring additional non-member data, i.e. ReCaLL [101], Probe (real) [62] and CON-ReCall [97], for the results in Table 4 as we use all non-member data for evaluation. Generating additional, synthetic non-member data for the trap sequences in Table 5 was feasible by following code made available by Meeus et al. [70].
Finally, for the results in tables 5 and 4, we compute the AUC on $1,000$ bootstrapped subsets of members and non-members and report both the mean and standard deviation across all results [5].
Appendix B Results for document-level MIAs
@@
MIA methodology. We instantiate sequence- and document-level MIAs from the literature. We report both the AUC achieved for all sequences ($50,000$ in total) as well as the AUC when the sequence-level membership scores are scaled up to the document-level. For this, we consider the method proposed by Shi et al. [86] and described in Sec. 5.2. Namely, a membership inference signal is computed on the $k=25$ sequences $s_{tj}$ of $200$ words per document $d_{t}$. Next, an optimal threshold $\tau_{\text{seq}}$ is determined for the inference signal to maximize accuracy on the sequence-level MIA. By classifying the selected sequences using the optimal threshold, the average binary prediction is then computed for a document, and is used as document-level scoring function $y_{\text{doc}}(\mathcal{M},d_{t})$, or $y_{\text{doc}}(\mathcal{M},d_{t})=\frac{1}{k}\sum_{j=1}^{k}\hat{y}_{\text{seq}}(\mathcal{M},s_{tj})$. Again, we do not consider the MIAs requiring additional non-member data, i.e. ReCaLL [101], Probe (real) [62] and CON-ReCall [97], and leave out SaMIA [45] and Neighborhood [68] due to the large computational costs required for this number of sequences.
@@
Table 8: Benchmarking sequence and document-level MIAs. Left: MIA AUC for ArXiv papers ($25$ sequences of $200$ words for both $1,000$ members and non-members) randomly sampled from train and test split from The Pile [2], attacking Pythia-dedup-6.9B [7]. Right: MIA AUC for ArXiv papers (same length) randomly sampled from the RDD setup, attacking OpenLLaMA-7B [32].

### v3

@@ VI-D Post-hoc control methods
Beating the model-less baseline. A potential solution is to revisit the baseline used to infer binary membership. MIAs against ML models traditionally use a 50/50 split of membership for evaluation, resulting in a random guess AUC of $0.5$ [87, 11, 18, 56, 83, 95, 35, 52, 74, 81, 105]. Alternatively, Jayaraman et al. [42] argues that a balanced prior for membership is not realistic in practice and proposes to evaluate MIAs with skewed priors. More recently, Kazmi et al. [47] proposes privacy auditing of ML models using a baseline from a (model-less) classifier. Further, Chang et al. [16] evaluate a bag of words classifier on the MIMIR dataset [27] and use this as a baseline for their new MIA.
Hence, we could set the performance of the bag-of-words classifier as a prior. An effective MIA methodology should then outperform this baseline by leveraging access to the target model exploiting its memorization of members.
This solution has significant limitations in practice. First, the performances of the bag of words classifier applied to the widely used datasets collected post-hoc are near-perfect (Table II). With such a significant distribution shift between members and non-members, it becomes practically impossible for novel MIAs to surpass this baseline and might lead us to (incorrectly) conclude that the target model does not memorize. Second, such model-less baselines will always be a lower bound. In Sec. IV, we only used a simple bag-of-words classifier, since its performance was already substantial. Had this classifier performed closer to random guessing, a subtler distribution shift might still exist, potentially detectable by more sophisticated classifiers [111, 24].
TABLE VI: Recommended benchmarks to evaluate new MIAs against LLMs proposed in the future, highlighting benefits and limitations.
@@
Sequence The randomized split of The Pile [31], with additional deduplication, as made available by Duan et al. [27] in the MIMIR dataset [^2] https://huggingface.co/datasets/iamgroot42/mimir. We recommend using the split where sequences with more than 80% of $13$-gram overlapping are not removed, not considering data from Github. Suite of Pythia [6] and GPT-NEO [7] models. Allows for randomized MIA evaluation. MIAs are more challenging due to the inherent overlap between members and non-members This however represents a more realistic setup in general.
@@
Sequence The randomized unique traps sequences as made available by Meeus et al. [69], across all number of repetitions $(10,100,1000)$. CroissantLLM [30] Allows for randomized MIA evaluation in a real-world model, limiting the overlap between members and non-members, and across different memorization regimes. MIA performance might not be realistic, and might be specific to sequence characteristics such as length, and perplexity.
@@
Document Randomized set of ArXiv papers sampled from the train and test splits from The Pile [31], both available as complete documents as well as in sequences. Suite of Pythia [6] and GPT-NEO [7] models. Allows for randomized MIA evaluation. MIAs are more challenging due to the inherent overlap between members and non-members This however represent a more realistic setup in general.
@@
Document Set of ArXiv papers sampled close to the training data cut-off date for OpenLLaMA[32]. Suite of OpenLLaMA[32] models. Allows for MIA evaluation on data collected post-hoc with any distribution shift between members and non-members largely mitigated (RDD). Might still be subject to a distribution shift, and faces potential data overlap.
@@
Regression Discontinuity Design (RDD). In the domain of causal inference, Regression Discontinuity Design (RDD) is a technique used to study the effect of a treatment when the assignment to the treatment has not been randomized but is instead based on a cutoff point. By sampling observations in a principled manner close to this cutoff, RDD enables the estimation of the treatment effect even in the absence of a randomized setup. RDD is for instance widely used in econometrics [50], where cutoff points might include income levels for welfare programs or geographical borders when studying different policy implications.
@@
We instantiate this RDD setup on OpenLLaMA [32], a model for which the entire training dataset is known. As members, we consider papers from a month before the training data cutoff date (February ’23). For non-members, we consider papers released shortly after (March ’23).
@@
Table VIII (Appendix B) summarizes the AUC for both sequence- as well as document-level MIAs across the literature applied to this setup. We find that the performance of MIAs remains close to a random guess baseline, as also found in the case of a randomized test split (Sec. VI-A).
@@
While the RDD method we propose seems to offer a viable solution for the setup at hand (using OpenLLaMa [32] and ArXiv papers), its practical implementation may face limitations. First, it assumes knowledge of the training data cut-off date (which is likely quite different from the model release date) and relies on the availability of sufficient data released close to this date. For ArXiv papers, more than $10,000$ new papers are added every month [^3] https://arxiv.org/stats/monthly_submissions, allowing for ample data collection close to any cut-off date. While similar arguments can be made for other frequently updated data sources (e.g. news articles), the RDD approach could prove challenging for less frequently updated sources (e.g. crime novels). Second, the RDD method aims to significantly mitigate distribution shifts between members and non-members, but its quantification faces similar issues as discussed above. Lastly, the RDD setup does not address the issue of potential overlap between members and non-members. This might be preferred if one wants to calibrate realistic MIA performance in practice, but maybe less so when benchmarking new methods.
@@
Debiasing techniques. Eichler et al.[29] introduced two novel methods to mitigate bias between members and non-members collected post-hoc. The first method creates ‘non-biased’ datasets, where members and non-members share similar n-gram distributions. The second one creates ‘non-classifiable’ datasets, which resist model-less baselines like those used here and by Das et al.[21]. While authors show these techniques to eliminate biases present in the original datasets collected post-hoc, this approach faces similar limitations as other post-hoc control methods. Fully eliminating known biases, not introducing new biases and accurately evaluating their presence, remains a practical challenge.
VII Comparison to concurrent work
Concurrently with this work (the preprints were released within the same week), Das et al. [21] propose to use bag of words classifiers, which they call a ‘blind’ baseline, to evaluate MIAs against LLMs. They show bag of word classifiers to outperform MIAs in two datasets: Shi et al. [85] and Meeus et al. [68], discovering significant distribution shifts. As a potential solution, they propose the use of a randomized test split as made available for some LLMs.
We independently introduce a bag of words classifier, showed them to outperform MIAs, and identified distribution shifts in $6$ datasets [68, 85, 56, 104], including the three used by Das et al. [21]. We further analyzed how post-hoc data deduplication such as employed to create the MIMIR dataset [27] might introduce new biases, and identify a range of recent work on MIAs against LLMs evaluated in flawed setups (Table I). Lastly, we carefully examine different potential solutions going forward (Sec. VI), evaluate a wide range of MIAs in these setups (Tables IV and V) and propose a recommended set of benchmarks (Table VI).
VIII Conclusion
@@
We identified that recent research on MIAs against LLMs often uses flawed evaluation setups. In the last year, research has proposed at least $10$ different MIAs, whose performance has been impossible to fairly compare and benchmark.
@@
After carefully examining potential solutions, we identify a set of benchmarks to use when new MIAs are proposed and summarize our findings in Table VI. For sequence-level MIAs, we propose a combination of the MIMIR dataset [27] (with limited deduplication) and the randomized trap sequences provided by Meeus et al. [69]. MIMIR enables a randomized setup for MIA evaluation across different model sizes [6, 7] and, despite its overlap issues, serves as a realistic benchmark for MIAs in practice. The trap sequences allow for benchmarking new MIA methodologies across different levels of memorization regimes; however, the corresponding results may depend on the specific properties of these sequences.
@@
Finally, we encourage future research to always include a model-less baseline and to remain mindful of the inherent challenges (e.g. overlap, scaling from finetuning to pretraining) before interpreting results.
Further information and resources to implement these benchmarks, alongside all the MIAs considered in this work, can be found in the Github repository we make available.
Appendix A1 Detailed MIA implementation
A. Detailed MIA implementation
Table VII elaborates on the implementation details used for the MIAs from the literature we consider. Note that some MIAs, i.e. Loss [105], Zlib [15] and Lower [15], do not depend on the choice of certain hyperparameters.
TABLE VII: MIA Implementation Details Summary
@@ A. Detailed MIA implementation
PAC [104] Using the implementation provided by Ye et al. [104], we generate $10$ augmentations with alpha parameter of $0.3$. We the compute polarized distance using $k_{\text{max}}=0.05$ and $k_{\text{min}}=0.3$ following the best performing results reported in the original paper.
@@
Note that we do not consider the MIAs requiring additional non-member data, i.e. ReCaLL [100], Probe (real) [60] and CON-ReCall [96], for the results in Table IV as we use all non-member data for evaluation. Generating additional, synthetic non-member data for the trap sequences in Table V was feasible by following code made available by Meeus et al. [69].
Finally, for the results in tables V and IV, we compute the AUC on $1,000$ bootstrapped subsets of members and non-members and report both the mean and standard deviation across all results [4].
B. Results for document-level MIAs
@@
MIA methodology. We instantiate sequence- and document-level MIAs from the literature. We report both the AUC achieved for all sequences ($50,000$ in total) as well as the AUC when the sequence-level membership scores are scaled up to the document-level. For this, we consider the method proposed by Shi et al. [85] and described in Sec. V-B. Namely, a membership inference signal is computed on the $k=25$ sequences $s_{tj}$ of $200$ words per document $d_{t}$. Next, an optimal threshold $\tau_{\text{seq}}$ is determined for the inference signal to maximize accuracy on the sequence-level MIA. By classifying the selected sequences using the optimal threshold, the average binary prediction is then computed for a document, and is used as document-level scoring function $y_{\text{doc}}(\mathcal{M},d_{t})$, or $y_{\text{doc}}(\mathcal{M},d_{t})=\frac{1}{k}\sum_{j=1}^{k}\hat{y}_{\text{seq}}(\mathcal{M},s_{tj})$. Again, we do not consider the MIAs requiring additional non-member data, i.e. ReCaLL [100], Probe (real) [60] and CON-ReCall [96], and leave out SaMIA [45] and Neighborhood [66] due to the large computational costs required for this number of sequences.
@@
TABLE VIII: Benchmarking sequence and document-level MIAs. Left: MIA AUC for ArXiv papers ($25$ sequences of $200$ words for both $1,000$ members and non-members) randomly sampled from train and test split from The Pile [1], attacking Pythia-dedup-6.9B [6]. Right: MIA AUC for ArXiv papers (same length) randomly sampled from the RDD setup, attacking OpenLLaMA-7B [32].